OpenStudy (anonymous):

i need the equation of the line that satisfies the following conditions: it passes through the point of intersection of the curves y=x^2-8x+8 and 2+x=y+2 and it is also perpendicular to the equation 2y-3+6x=0

OpenStudy (anonymous):

I'm going to try and help for this one give me a sec

OpenStudy (anonymous):

FOR NOW YOUR SLOPE IS THE NEGATIVE RECIPROCAL OF THE EQUATION 2Y-3+6X=0 WHICH IS 1/3 BECAUSE SLOPE OF THE EQUATION ABOVE IS -3 WHEN RE-WRITTEN.

OpenStudy (anonymous):

so if i find the intersection by substituting y=x^2-8x+8 to 2 +x=y+2? and then use those points to find my equation?

OpenStudy (anonymous):

RIGHT, THE OTHER TWO EQUATIONS INTERSECT AT (1,1) AKA (X1,Y1) AND THE EQUATION FOR A LINE THAT INTERSECTS A POINT IS Y-Y1=M(X-X1)

OpenStudy (anonymous):

but what about the other point

OpenStudy (anonymous):

ive got to intersections

OpenStudy (anonymous):

RIGHT! (8,8) YOU CAN USE EITHER POINT.. POINT INTERCEPT FORM OF AN EQUATION ONLY REQUIRES ONE SET OF POINTS AND THE SLOPE M=1/3.. SO HOPE I HELPED A BIT

OpenStudy (anonymous):

I MEAN POINT SLOPE HAHA

OpenStudy (anonymous):

ok ok thanks man

OpenStudy (anonymous):

i got it from here

OpenStudy (anonymous):

RIGHT ON! GOOD LUCK

OpenStudy (anonymous):

y=x/3+2/3 or y=x/3+16/3